OpenStudy (anonymous):

A 3-mi cab ride costs $3.00.A 6-mi cab ride cost $4.80.Find a linear equation that models costs c as a function of distance d.

OpenStudy (sriram):

y=mx +c where y is cost x is dist m and c const now 3= 3m +c n 4.8= 6m+c solve them to get m=0.6 and c=1.2

OpenStudy (anonymous):

so it would be c=0.06d+1.20

OpenStudy (sriram):

meaning for every mile the rateincreases by 0.6$ (hence it increased from 3 to 4.8 for 3 miles increase and the cost at the beginnin is 1.2 $